# Towards Long-Term User Welfare in Recommender Systems via Creator-Oriented Information Revelation

### v1

@@ Abstract.
Improving the long-term user welfare (e.g., sustained user engagement) has become a central objective of recommender systems (RS). In real-world platforms, the creation behaviors of content creators plays a crucial role in shaping long-term welfare beyond short-term recommendation accuracy, making the effective steering of creator behavior essential to foster a healthier RS ecosystem. Existing works typically rely on re-ranking algorithms that heuristically adjust item exposure to steer creators’ behavior. However, when embedded within recommendation pipelines, such a strategy often conflicts with the short-term objective of improving recommendation accuracy, leading to performance degradation and suboptimal long-term welfare.
The well-established economics studies offer us valuable insights for an alternative approach without relying on recommendation algorithmic design: revealing information from an information-rich party (sender) to a less-informed party (receiver) can effectively change the receiver’s beliefs and steer their behavior. Inspired by this idea, we propose an information-revealing framework, named Long-term Welfare Optimization via Information Revelation (LoRe). In this framework, we utilize a classical information revelation method (i.e., Bayesian persuasion) to map the stakeholders in RS, treating the platform as the sender and creators as the receivers. To address the challenge posed by the unrealistic assumption of traditional economic methods, we formulate the process of information revelation as a Markov Decision Process (MDP) and propose a learning algorithm trained and inferred in environments with boundedly rational creators. Extensive experiments on two real-world RS datasets and across different creator modeling approaches demonstrate that our method can effectively outperform existing fair re-ranking methods and information revealing strategies in improving long-term user welfare.
@@ 1. Introduction
With the rapid growth of user-generated content (UGC) on the Web (Krumm et al., 2008), modern recommender platforms (e.g., TikTok, YouTube) have evolved into complex ecosystems where users and content creators continuously interact and influence each other (Bobadilla et al., 2013; Ye et al., 2025). In such dynamic environments, platforms are now shifting their focus from solely optimizing short-term objectives, such as maximizing users’ immediate engagement behaviors (e.g., clicks or purchases), to fostering long-term user welfare (i.e., sustained user engagement), aiming to ensure the lasting health and sustainability of the ecosystem (Xu et al., 2023; Guo et al., 2023).
@@
To enhance the long-term user welfare of recommender systems (RS), previous works have shown that the creators’ creation behavior (i.e., creating new items) is a key factor (Ye et al., 2025; Xu et al., 2023; Yao et al., 2024) beyond recommendation accuracy (i.e., prediction accuracy of user behavior). This is because creators continuously upload new items based on previous user feedback, which shapes the item pool of RS and indirectly influences what users are exposed to in the long run (Ben-Porat and Tennenholtz, 2018; Ben-Porat et al., 2019; Boutilier et al., 2024). Considering the importance of creators’ behavior, most existing works utilize heuristic re-ranking algorithms to steer creators’ content creation. For example, they increase item exposure for small creators to enhance creator retention and stimulate content creation (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024). However, previous works have shown that such a re-ranking strategy, when embedded within recommendation pipelines, often conflicts with the short-term objectives of improving accuracy, inevitably trading off recommendation performance and harming users’ short-term experience (Ye et al., 2024; Xu et al., 2023). Therefore, how to steer creators’ content creation behavior beyond relying on recommendation algorithmic design to enhance the long-term welfare remains a significant challenge in RS.
@@
The investigation of behavioral steering among strategic agents has a long history in economics (Milgrom and Roberts, 1987; Dierkens, 1991; Mishra et al., 1998; Crawford et al., 2018), offering valuable insights for enhancing the long-term welfare of RS. In economics studies, a common approach is to leverage information revelation under conditions of information asymmetry (i.e., agents possess unequal amounts of information) to influence agents’ behavior. Such a strategy aims to design an information revelation strategy, which specifies what information the information-rich party (sender) should disclose to the less-informed party (receiver). By revealing certain information unavailable to the receiver, the receiver can update their beliefs and adjust their strategy accordingly, leading to observable behavioral changes.
@@
Inspired by this idea, we can steer creators’ behavior to enhance long-term welfare by revealing information, viewing the platform as the sender and creators as the receivers. Such a strategy is feasible and practical in real-world platforms, because information asymmetry naturally exists between the platform and creators in RS (Prasad et al., 2023; Boutilier et al., 2024; Ye et al., 2025). Similar but simpler idea have been adopted in industrial practice, offering real-world scenarios and potential applications for more advanced information revelation approaches. For example, platforms like TikTok provide creators with suggestions to produce trending content (see Appendix A).
@@
However, directly applying information revelation to RS is challenging. Traditional methods in economics studies (Yao et al., 2024; Prasad et al., 2023; Boutilier et al., 2024) typically assume that receivers are completely rational, i.e., they can always select the optimal strategy to maximize their utility given the available information. In contrast, real-world creators are bounded rational (Simon, 1990; Kahneman and Tversky, 2013), because their decision-making is constrained by the limited cognitive capacity, attention, and information-processing ability of humans.
@@
To address this challenge, we propose our information-revealing framework tailored for RS, named Long-term Welfare Optimization via Information Revelation (LoRe). Specifically, we map the key components of classical information revelation methods in economics (i.e., Bayesian persuasion (Gan et al., 2022a; Kamenica and Gentzkow, 2011)) onto the corresponding stakeholders in RS. Based on this mapping, we formulate the platform’s information revelation process as a Markov Decision Process (MDP) (Puterman, 1990) and propose a learning algorithm to optimize the information revelation strategy. Instead of relying on the assumption of completely rational agents, LoRe is trained and inferred within a realistic reinforcement learning (RL) environment, where bounded rational users and creators interact with each other.
@@
Unlike previous works that rely on heuristic exposure allocation in the re-ranking stage to steer creators’ behavior (Guo et al., 2023), LoRe serves as an independent module for the platform and does not depend on any specific recommendation algorithm. Experiments on two real-world recommendation datasets demonstrate that LoRe outperforms both creator-fair re-ranking methods and simple information-revealing strategies in improving long-term user welfare (i.e., sustained user clicks). Moreover, since LoRe operates independently, it can be seamlessly integrated with existing ranking or re-ranking algorithms to further enhance long-term welfare.
@@
(1) We highlight that steering creators’ behavior through information revealing provides an effective approach to enhancing long-term user welfare beyond relying on recommendation algorithmic design.
(2) Inspired by economic studies, we propose a novel information revelation method, LoRe, which models the revelation process as a Bayesian persuasion problem and learns to optimize the revelation strategy in environments where boundedly rational agents interact.
(3) Extensive experiments on real-world datasets demonstrate that our method can effectively improve the long-term user welfare of RS by steering creators’ content creation.
@@ 2. Related Work
Multi-Stakeholder RS. Multi-stakeholder RS have received significant attention in recent years in both academic research and industrial practice (Abdollahpouri et al., 2020; Chen et al., 2025). Unlike traditional RS, which primarily focuses on providing the most suitable items based on each user’s request (He et al., 2017; Rendle et al., 2012), multi-stakeholder RS also takes into account the interests of platforms and creators, aligning with the long-term goals of many multi-stakeholder real-world platforms (e.g., TikTok, YouTube). Most existing studies on multi-stakeholder RS can be broadly categorized into two directions: (1) designing better RS algorithms that aim to achieve multi-stakeholder objectives (e.g., provider fairness (Ye et al., 2024; Xu et al., 2025), platform profit (Pei et al., 2019; Azaria et al., 2013)) while maintaining recommendation accuracy; (2) incorporating creators and platforms into the RS environment to enable more comprehensive evaluation of multi-stakeholder RS. For example, some studies focused on developing re-ranking algorithms that allocate exposure resources fairly among different providers to achieve different provider fairness objectives (e.g., Gini coefficient (Do and Usunier, 2022), Max-Min Fairness (Ye et al., 2024; Xu et al., 2023)). Ye et al. (2025) introduces an LLM-based creator simulator, CreAgent, aiming to achieve a more realistic evaluation of multi-stakeholder RS. However, existing works mainly focus on developing and evaluating RS algorithms, and do not consider another crucial side, i.e., steering creators via creation suggestions to indirectly optimize the long-term welfare of RS.
@@
Information asymmetry in RS. A relevant topic to our work is the literature considering the information asymmetry in RS (Prasad et al., 2023; Yao et al., 2023, 2024; Boutilier et al., 2024; Ye et al., 2025). Boutilier et al. (2024); Ye et al. (2025) highlight the significant information asymmetry in RS, but they do not proposed a method to improve the welfare of RS under such asymmetry. Prasad et al. (2023) propose a prompting policy that promises to provide creators a certain level of user clicks to influence the their content creation. Yao et al. (2023, 2024) suggest influencing content creation by adjusting reward weights to encourage different types of content. However, these studies model creators as completely rational agents, ignoring the bounded rationality observed in real-world creators, which limits the applicability and performance of their methods.
Information revelation under information asymmetry. Our work is also related to the literature on information revelation under information asymmetry in economics, where a prominent approach is Bayesian persuasion (Kamenica and Gentzkow, 2011). In Bayesian persuasion, an informed party (sender) designs an information revelation strategy to reveal the optimal information to a less informed party (receiver) (Kamenica and Gentzkow, 2011). The most relevant Bayesian persuasion models to our work consider one sender (the platform) and multiple receivers (the creators) (Wang, 2013; Arieli and Babichenko, 2019; Castiglioni et al., 2021; Bernasconi et al., 2023; Zhao et al., 2025), where the sender simultaneously reveals information to multiple receivers. These studies investigate how to design information disclosure strategies to maximize the sender’s utility in a static setting. There is also a stream of research on dynamic Bayesian persuasion (Gan et al., 2022b, a; Wu et al., 2022), which examines repeated interactions between a sender and a receiver. However, because of their strong assumption of completely rational receivers, these models cannot be directly applied to real-world RS scenarios, where creators typically exhibit bounded rationality.
@@ 3. Preliminaries
In this section, we first introduce the settings of real-world RS. Then we introduce the basic idea of Bayesian Persuasion. Finally, we describe how to model creators as bounded rational agents.
@@ 3.1. Recommender Systems
Interactions between different stakeholders. We model the interaction between the platform, creators, and users as a dynamic process with total $T$ interaction rounds. At the beginning of each round $t\in\{1,2,\dots,T\}$, each creator $c_{j}$ will produce an item with the genre in $\mathcal{G}$. The items produced by creators in each round will be added to $\mathcal{I}$. Then, some users will arrive and access the platform. For each accessing user, the platform needs to select $k$ items from $\mathcal{I}$ for recommendation. Each user can choose to click or skip the items recommended to him. After that, the platform records the interaction between users and items. Specifically, we use $\bm{Y}$ to represent the click matrix, where $\bm{Y}_{i,j,t}=1$ if item $i\in\mathcal{I}$ is clicked by user $u_{j}$ in round $t$ and 0 otherwise. In addition, we use $\mathcal{H}_{c_{j}}=\{(i,k_{i})|i\in\mathcal{I}\}$ to denote historical information of items created by creator $c_{j}$, where $(i,k_{i})\in\mathcal{H}_{c_{j}}$ means that the item $i\in\mathcal{I}$ produced by creator $c_{j}$ has been clicked $k_{i}$ times. For easy notation, let $\mathcal{H}=\{\mathcal{H}_{c_{1}},\mathcal{H}_{c_{2}},\dots,\mathcal{H}_{c_{|\mathcal{C}|}}\}$ denote the set of all creators’ historical information.
@@
Information asymmetry in RS. In real-world RS, the platform has the historical information $\mathcal{H}$ of all creators and the interaction information $\bm{Y}$ of all users and items. In contrast, each creator $c_{j}$ only has its own historical information $\mathcal{H}_{c_{j}}$. Obviously, there is a significant information asymmetry between platform and creators. Therefore, how to improve the long-term welfare under information asymmetry is an unexplored and urgent problem to be addressed.
Figure 1. The overall framework of our method LoRe, which consists two phases: Online Inference and Offline Training.
3.2. Bayesian Persuasion
@@
Information asymmetry formulation. The standard Bayesian persuasion considers an idealized model with a sender and a receiver. Both the receiver’s utility function $\mu_{r}(a_{r},w)$ and the sender’s utility function $\mu_{s}(a_{r},w)$ depend on receiver’s action $a_{r}\in\mathcal{A}_{r}$ and state $w\in\mathcal{W}$. The state $w$ follows a cumulative distribution function $F(w)$ with probability density function $f(w)$. It is assumed that both sender and receiver know $F(w)$, but the sender exactly knows the real state $w$ after $w$ is sampled. Thus, the sender have more information than the receiver, which causes the information asymmetry.
Bayesian persuasion under information asymmetry. The receiver will choose an optimal action $a_{r}^{*}$ to maximize her expected utility, and the sender will also obtain a corresponding utility. However, the sender can use a so-called signaling scheme to reveal the receiver some information, and then influence receiver’s action. Specifically, we use $\sigma=(\mathcal{A}_{s},\pi)$ to denote a signaling scheme, where $\mathcal{A}_{s}$ is the set of possible signals (messages) and $\pi:\mathcal{W}\times\mathcal{A}_{s}\mapsto[0,1]$ is the signaling strategy which maps the state to a distribution of signals. For instance, $\pi(a_{s}|w)$ denotes the probability of sending signal $a_{s}\in\mathcal{A}_{s}$ when the real state is $w$. According to the revelation principle (Myerson, 1979), each signal $a_{s}$ can be viewed as an action suggestion for receiver, that is $|\mathcal{A}_{r}|=|\mathcal{A}_{s}|$, where $a_{s}\in\mathcal{A}_{s}$ means suggesting receiver taking corresponding action $a_{r}\in\mathcal{A}_{r}$ and receiver will follows sender’s suggestion [^1] Some constraints can be used to ensure that the receiver will follow the sender’s suggestion, and these constraints do not affect the optimal solution. For details, please refer to the reference (Kamenica and Gentzkow, 2011).. By designing the optimal strategy $\pi^{*}$, the sender can improve both the utility of sender and receiver under information asymmetry (Kamenica and Gentzkow, 2011).
@@
Table 1. Correspondence between elements in the Bayesian Persuasion and RS. See Section 4 for undefined notations.
Bayesian Persuasion Recommender System
@@
state $w$ content distribution $\bm{e}$ and trust $\bm{d}$
signal set $\mathcal{A}_{s}$ suggestion set $\mathcal{A}_{p}$
signaling strategy $\pi$ information revealing strategy $\pi$
3.3. Bounded Rationality Modeling
Motivated by the advantage of Bayesian persuasion under information asymmetry, we adopt the idea of this framework to guide how the platform reveals information in RS. Simply put, we allow the platform to send suggestions to creators, i.e., suggesting creators creating specific genres of content, which is called the platform’s information revelation strategy.
In the standard Bayesian persuasion, receivers are assumed to be completely rational, which means they can precisely choose an action $a_{r}^{*}$ to maximize their utility $\mu_{r}(a_{r},w)$. This also allows senders to design signaling strategy that ensure that receivers completely follow the sender’s action suggestion. However, creators in reality are typically bounded rational, as they cannot compute a specific utility function to maximize their utility. Thus, platform in RS cannot make sure that creators will follow platform’s suggestion. To model creators’ bounded rationality, inspired by Prasad et al. (2023), we design a trust-based model, where each creator $c_{i}$ has a trust degree $d_{{}_{i}}\in[0,1]$. When creator $c_{i}$ receives platform’s suggestion, the probability that creator $c_{i}$ follows the suggestion is $d_{{}_{i}}$ .
If creator $c_{i}$ dose not follow platform’s suggestion, we assume that he can use any strategy that does not perform utility calculation to choose an action, such as randomly selecting from the genres of previously created content, or selecting the most popular genre of previously created content, etc. That is, we do not restrict the creators’ decision model, and any decision model can be embedded in our approach that is introduced in next section. In the experiment section, to verify the universality of our method for different creator decision models, we use different model to simulate creators’ decision-making, such as LLM-based simulation model and heuristic-based simulation models.
We correspond the elements of RS with those in the Bayesian persuasion, as shown in Table 1.
$\bullet$ The platform and creators are regarded as the sender and receivers, respectively. This is because the platform, who possesses more information, can guide the creator’s creation by revealing information.
$\bullet$ The content distribution in RS and creators’ trust degree are viewed as the state in Bayesian persuasion.
$\bullet$ The suggestion that platform sends to the creators is considered as the signal in Bayesian persuasion.
$\bullet$ The platform’s strategy of sending suggestions (information revelation strategy $\pi$) is corresponds to the signaling strategy $\pi$ in Bayesian persuasion.

### v2

# Enhancing Long-Term Welfare in Recommender Systems: An Information Revelation Approach

@@ Abstract.
Improving long-term user welfare (e.g., sustained user engagement) has become a central objective in recommender systems (RS). In real-world platforms, the creation behaviors of content creators play a crucial role in shaping long-term welfare beyond short-term recommendation accuracy, making effective steering of creator behavior essential for fostering a healthy RS ecosystem. Existing approaches typically rely on re-ranking algorithms that heuristically adjust item exposure to influence creators’ behavior. However, when integrated into recommendation pipelines, such methods often conflict with the short-term objective of improving recommendation accuracy, leading to performance degradation and suboptimal welfare.
The well-established studies in economics provide valuable insights into an alternative approach without relying on recommendation algorithm design: revealing information from an information-rich party (sender) to a less-informed party (receiver) can effectively influence the receiver’s beliefs and steer their behavior. Inspired by this idea, we propose an information-revealing framework, named Long-term Welfare Optimization via Information Revelation (LoRe). In LoRe, we adopt a classical information revelation method (i.e., cheap talk) to map the stakeholders in RS, treating the platform as the sender and creators as the receivers. To address the challenge posed by the unrealistic assumption of traditional economic method, we formulate the process of information revelation as a Markov Decision Process (MDP) and develop a learning algorithm to optimize the information revelation strategy. Extensive experiments on two real-world RS datasets and across different creator modeling approaches demonstrate that our method can effectively outperform existing re-ranking methods and reward-based creator-influencing method in improving long-term user welfare.
@@ 1. Introduction
With the rapid growth of user-generated content (UGC) (Krumm et al., 2008), modern recommender platforms (e.g., TikTok, YouTube) have evolved into complex ecosystems where users and content creators continuously interact and influence each other (Bobadilla et al., 2013; Ye et al., 2025). In such dynamic environments, platforms are now shifting their focus from solely optimizing short-term objectives, such as maximizing users’ immediate engagement behaviors (e.g., clicks), to fostering long-term user welfare (i.e., sustained user engagement), aiming to ensure the lasting health and sustainability of the platform ecosystem (Xu et al., 2023; Guo et al., 2023).
@@
To enhance the long-term user welfare of recommender systems (RS), previous studies have shown that the creators’ creation behavior (i.e., creating new items) is a key factor (Ye et al., 2025; Xu et al., 2023; Yao et al., 2024) beyond recommendation accuracy (i.e., prediction accuracy of user behavior). This is because creators continuously upload new items based on previous user feedback, thereby shaping the item pool of RS and indirectly influences what users are exposed to in the long run (Ben-Porat and Tennenholtz, 2018; Ben-Porat et al., 2019; Boutilier et al., 2024). Given the importance of creator behavior, most existing works utilize heuristic re-ranking algorithms to steer creators’ content creation. For example, they increase item exposure for smaller creators to enhance creator retention and stimulate content creation (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024). However, prior work have shown that such re-ranking strategies, when embedded within recommendation pipelines, often conflict with the short-term objectives of improving accuracy, inevitably trading off recommendation performance and harming users’ short-term experience (Ye et al., 2024; Xu et al., 2023). Consequently, how to steer creators’ content creation behavior without relying on recommendation algorithmic design to enhance the long-term welfare remains a significant challenge in RS.
@@
The investigation of behavioral steering among strategic agents has a long history in economics (Milgrom and Roberts, 1987; Dierkens, 1991; Mishra et al., 1998; Crawford et al., 2018), offering valuable insights for enhancing the long-term welfare of RS. In economics studies, a common approach is to leverage information revelation under the situation of information asymmetry (i.e., agents possess unequal amounts of information) to influence agents’ behavior. Such a strategy focuses on designing an information revelation strategy that determines what information an information-rich party (sender) should disclose to a less-informed party (receiver). By revealing certain information unavailable to the receiver, the receiver can update the beliefs and adjust the action accordingly, resulting in observable behavioral changes. The revealed information in such situation can be generally interpreted as a suggestion that guides the receiver’s action.
@@
Inspired by this idea, we can steer creators’ behavior to enhance long-term welfare by revealing information, viewing the platform as the sender and creators as the receivers. Such a strategy is feasible and practical in real-world platforms, because information asymmetry naturally exists between the platform and creators in RS (Prasad et al., 2023; Boutilier et al., 2024; Ye et al., 2025). Similar but simpler idea have been adopted in industrial practice, offering real-world scenarios and potential applications for more advanced information revelation approaches. For instance, platforms like TikTok and YouTube provide creators with suggestions to produce trending content.
@@
However, directly applying information revelation to RS is challenging. Traditional economic approaches (Yao et al., 2024; Prasad et al., 2023; Boutilier et al., 2024) typically assume that receivers are completely rational, i.e., they always select the optimal strategy to maximize their utility. Under this assumption, receivers’ behavior can be computed explicitly, enabling the sender to optimally send information to maximize its own utility. In contrast, real-world creators are boundedly rational (Simon, 1990; Kahneman and Tversky, 2013), because their decision-making is constrained by the limited cognitive capacity, attention, and information-processing ability of humans. Consequently, it is difficult to calculate creators’ behavior in real-world precisely, preventing the platform from directly computing the optimal information to reveal.
@@
To address this challenge, we propose an information revelation framework tailored for RS, named Long-term Welfare Optimization via Information Revelation (LoRe). Specifically, we map the key components of classical information revelation method in economics studies (i.e., cheap talk model (Farrell and Rabin, 1996)) onto the corresponding stakeholders in RS. Based on this mapping, we formulate the platform’s information revelation process as a Markov Decision Process (MDP) (Puterman, 1990) and develop a learning algorithm to optimize the platform’s information revelation strategy.
@@
Unlike previous works that rely on heuristic exposure allocation in the re-ranking stage to steer creators’ behavior (Guo et al., 2023), LoRe serves as an independent module for the platform and does not depend on any specific recommendation algorithm. Experiments on two real-world recommendation datasets demonstrate that LoRe outperforms both creator-fair re-ranking methods and reward-based creator-influencing method in improving long-term user welfare (i.e., sustained user clicks). Moreover, since LoRe operates independently, it can be seamlessly integrated with existing ranking or re-ranking algorithms to further enhance long-term welfare.
In our experiments, to more faithfully simulate the real-world RS environments, we adopt the multi-stakeholder simulation framework proposed by Ye et al. (2025), which leverages LLM-based agents to simulate the dynamic behaviors of users and creators. Specifically, these agents simulate user click-through behavior and creator content generation. This approach is grounded in recent research which demonstrate that agents empowered by LLMs can effectively emulate user preferences and the creation behavior of bounded rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025).
@@
(1) We highlight that influencing creators’ creation behavior through information revelation provides an effective approach to enhancing long-term user welfare beyond relying on recommendation algorithm design.
(2) Inspired by economic studies, we propose a novel information revelation method, LoRe, which models the revelation process as a MDP and learns to optimize the information revelation strategy in environments where boundedly rational agents interact.
(3) Extensive experiments on real-world datasets demonstrate that our method can effectively improve the long-term user welfare of RS by steering creators’ creation behavior.
@@ 2. Related Work
Multi-stakeholder RS. Multi-stakeholder RS have received significant attention in recent years in both academic research and industrial practice (Abdollahpouri et al., 2020; Chen et al., 2025). Unlike traditional RS, which primarily focuses on providing the most suitable items based on each user’s request (He et al., 2017; Rendle et al., 2012), multi-stakeholder RS also takes into account the interests of platforms and creators, aligning with the long-term goals of many multi-stakeholder real-world platforms (e.g., TikTok, YouTube). Most existing studies on multi-stakeholder RS can be broadly categorized into two directions: (1) designing better RS algorithms that aim to achieve multi-stakeholder objectives (e.g., provider fairness (Ye et al., 2024; Xu et al., 2025), platform profit (Pei et al., 2019; Azaria et al., 2013)) while maintaining recommendation accuracy; (2) incorporating creators and platforms into the RS environment to enable more comprehensive evaluation of multi-stakeholder RS. For example, some studies focused on developing re-ranking algorithms that allocate exposure resources fairly among different providers to achieve different provider fairness objectives (e.g., Gini coefficient (Do and Usunier, 2022), Max-Min Fairness (Ye et al., 2024; Xu et al., 2023)). Ye et al. (2025) introduces an LLM-based creator simulator, CreAgent, aiming to achieve a more realistic evaluation of multi-stakeholder RS. However, existing works mainly focus on developing and evaluating RS algorithms, and do not consider another crucial side, i.e., steering creators via information revelation to indirectly optimize the long-term welfare of RS.
@@
Information asymmetry in RS and information revelation. A relevant topic to our work is the literature considering the information asymmetry in RS (Prasad et al., 2023; Yao et al., 2023, 2024; Boutilier et al., 2024; Ye et al., 2025). Boutilier et al. (2024); Ye et al. (2025) highlight the significant information asymmetry in RS, but they do not proposed a method to address such a problem. Prasad et al. (2023) propose a prompting policy that promises to provide creators a certain level of user clicks to influence the their creation behavior. Yao et al. (2023, 2024) suggest influencing content creation by adjusting reward weights to encourage different types of content. However, these methods inevitably harm creators’ realized rewards and may introduce unfairness, as they directly alter the rewards that creators would otherwise receive. The literature on information revelation under information asymmetry in economics is also related to our work, where a prominent approach is cheap talk (Farrell and Rabin, 1996; Teacy, 2006; Chakraborty and Harbaugh, 2010; Ivanov, 2015). In cheap talk, an informed party (sender) designs an information revelation strategy to reveal the information to a less informed party (receiver). However, because of their strong assumption of completely rational receivers, these models cannot be directly applied to real-world RS scenarios where creators typically exhibit bounded rationality.
LLM-based simulator in RS. In recent years, given the superior semantic understanding and reasoning capabilities of LLMs, LLM-based simulators for recommender systems have attracted increasing attention and experienced rapid development. They are now widely regarded as a reliable technique for modeling user behavior in recommender systems (Ye et al., 2025; Ma et al., 2025; Zhu et al., 2025; Zhang et al., 2025). Initially, research (Ma et al., 2025; Zhu et al., 2025; Zhang et al., 2025) focused on simulating user behavior in RS, such as clicks, to obtain real-world user feedback for evaluating recommendation algorithms. Most recently, the field has progressed toward multi-stakeholder and ecosystem simulation, recognizing that RS operate within complex environments involving users, creators and platforms (Wang et al., 2024; Ye et al., 2025; Coppolillo et al., 2025). These studies demonstrated that LLM-based RS simulators can replicate the behaviors of both creators and users in real-world, providing a reliable framework to evaluate the performance of algorithms developed for such systems.
@@ 3.1. Recommender Systems
Interactions between different stakeholders. We model the interaction between the platform, creators, and users as a dynamic process with total $T$ interaction rounds. At the beginning of each round $t\in\{1,2,\dots,T\}$, each creator $c_{j}$ will produce an item with the genre in $\mathcal{G}$. The items produced by creators in each round will be added to $\mathcal{I}$. Then, some users will arrive and access the platform. For each accessing user, the platform needs to select $k$ items from $\mathcal{I}$ for recommendation. Each user can choose to click or skip the items recommended to him. After that, the platform records the interaction between users and items. Specifically, we use $\bm{Y}$ to represent the click matrix, where $\bm{Y}_{i,j,t}=1$ if item $i\in\mathcal{I}$ is clicked by user $u_{j}$ in round $t$ and 0 otherwise. In addition, we use $\mathcal{H}_{c_{j}}=\{(i,k_{i})\mid i\in\mathcal{I}\}$ to denote historical information of items created by creator $c_{j}$, where $(i,k_{i})\in\mathcal{H}_{c_{j}}$ means that the item $i\in\mathcal{I}$ produced by creator $c_{j}$ has been clicked $k_{i}$ times. For easy notation, let $\mathcal{H}=\{\mathcal{H}_{c_{1}},\mathcal{H}_{c_{2}},\dots,\mathcal{H}_{c_{|\mathcal{C}|}}\}$ denote the set of all creators’ historical information.
@@
Information asymmetry in RS. In real-world RS, the platform has the historical information $\mathcal{H}$ of all creators and the interaction information $\bm{Y}$ of all users and items. In contrast, each creator $c_{j}$ only has its own historical information $\mathcal{H}_{c_{j}}$. Obviously, there is a significant information asymmetry between platform and creators.
Figure 1. The overall framework of our method LoRe, which consists two phases: Inference and Training. During the inference phase, in each round, the platform sends suggestions to creators, who then produce content, after which users decide whether to click on the recommended items. The interaction data generated in each round are collected by the platform. In the training phase, the platform uses the data collected during inference to train the actor and critic networks.
3.2. Cheap Talk in Recommender Systems
@@
Cheap talk. The standard cheap talk model considers an idealized setting with a sender and a receiver. Both the receiver’s utility function $\mu_{r}(a_{r},w)$ and the sender’s utility function $\mu_{s}(a_{r},w)$ depend on receiver’s action $a_{r}\in\mathcal{A}_{r}$ and state $w\in\mathcal{W}$. In this model, only the sender exactly knows the true state $w$, whereas the receiver does not. Thus, the sender possesses more information than the receiver, giving rise to information asymmetry between sender and receiver.
When the receiver has no additional information, she chooses an optimal action $a_{r}^{*}$ to maximize her expected utility based on what she knows, and the sender correspondingly obtains utility $\mu_{s}(a_{r}^{*},w,b)$. However, the sender can influence the receiver’s action by revealing information (message). Specifically, we use $M$ to denote the set of possible messages the sender can send to receiver. The specific form of a message is immaterial, as long as it is commonly understood and interpretable by both parties. Intuitively, a specific message $m\in M$ can be viewed as a suggestion of the form: "Based on the available information, I recommend that you take action $\hat{a}_{r}$". Formally, we use $\pi:\mathcal{W}\times M\mapsto[0,1]$ to denote the sender’s information revelation strategy, where $\pi(m\mid w)$ represents the probability of sending message $m\in M$ when the real state is $w$. Upon receiving a message, the receiver updates her beliefs and chooses an optimal action accordingly. Previous researches (Farrell and Rabin, 1996; Teacy, 2006; Chakraborty and Harbaugh, 2010; Ivanov, 2015) shows that, by appropriately designing the information revelation strategy, the sender can improve the utilities of both the sender and the receiver under information asymmetry.
@@
Table 1. Correspondence between elements in Cheap Talk and RS. See section 4 for undefined notations.
Cheap Talk Recommender System
@@
state $w$ distribution of items $s$
message set $M$ suggestion set $\mathcal{A}_{p}$
information revelation strategy $\pi$ suggestion strategy $\pi_{\theta}$
Cheap talk in Recommender Systems. Motivated by the advantage of cheap talk under information asymmetry, we adopt this framework to guide how the platform reveals information in RS, where information asymmetry similarly exists between the platform and content creators. Simply put, we allow the platform to send suggestions to creators, such as "I suggest that you create an item of genre $g_{i}\in\mathcal{G}$, as this may attract more clicks or likes". We map the key components of RS to their counterparts in the cheap talk model, as summarized in Table 1. A detailed description of these correspondences is provided below.
We model the platform and the creators as the sender and the receivers, respectively. This modeling choice is motivated by the fact that the platform possesses richer information and can influence creators’ creation behavior through information revelation. Moreover, creators’ creation behavior shapes the item pool of RS, thereby indirectly determining the content to which users are exposed over time and, consequently, affecting long-term user welfare. Therefore, by steering creators’ creation behavior, platforms can indirectly improve long-term user welfare.
In the cheap talk model, the receiver’s utility depends on the actual state, which is unobserved by the receiver. Analogously, in RS, we represent the state as the distribution of item genres, since this distribution has a substantial impact on clicks, likes, and other performance metrics of creators’ items, and is typically unknown to individual creators. For example, if an increasing number of creators produce items of the same genre, competition within that genre intensifies, reducing the traffic received by each creator. We denote the set of possible states by $\mathcal{S}$.
As the platform is allowed to send suggestions to creators, the messages in the cheap talk model correspond to the platform’s suggestions in RS. We use $\mathcal{A}_{p}=\{a_{i}\}_{i=0}^{|\mathcal{G}|}$ to denote the set of platform’s possible suggestions, where $a_{0}$ represents sending no suggestion, and $a_{i}\ (i\neq 0)$ represents the suggestion: "I recommend that you create item with genre $g_{j}$".
Corresponding to the sender’s information revelation strategy in the cheap talk model, in RS, we use $\pi_{\theta}:\mathcal{S}\times\mathcal{A}_{p}\mapsto[0,1]$ to represent the platform’s suggestion strategy. Specifically, $\pi_{\theta}(a_{i}\mid s)$ is the probability that the platform sends suggestion $a_{i}$ when the state is $s\in\mathcal{S}$.
In the traditional cheap talk model, the sender can compute an optimal information revelation strategy to maximize his own utility because the receiver is assumed to be completely rational, and thus the receiver’s behavior in response to any message can be precisely inferred. In contrast, in RS, creators are not completely rational and their creation behavior after receiving platform’s suggestions is generally unknown to the platform. As a result, the platform cannot directly compute the optimal suggestion strategy. Consequently, the theoretical results of the classic cheap talk model cannot be directly applied to RS, making the optimization of the platform’s suggestion strategy a challenging problem.